OpenStudy (jjuden):

evaluate 6=sqrt 14+x when x=2

OpenStudy (jjuden):

\[6+\sqrt{14+x}\]

OpenStudy (triciaal):

6 + rt (14 +2) = 6 + rt(16) rt 16 = +/_ 4 6 + 4 : 6-4 10 ; 2